OpenStudy (anonymous):

[6.02] Use a system of equations to solve the following problem: Charles High School has a total of 94 boys and girls who play sports. If the number of boys is 22 more than twice the number of girls, how many boys play sports at this high school? 70 48 24 80

OpenStudy (jamesj):

No. You may not top post and ignore the answer that was being given to you.

OpenStudy (jamesj):

Like I said the last time you posted it, set up two variables B = number of boys G = number of girls What now are two equations you can write down with B and G?

OpenStudy (anonymous):

idk

OpenStudy (jamesj):

"Charles High School has a total of 94 boys and girls who play sports" How do you turn that into one equation involving B and G? Any ideas?

OpenStudy (anonymous):

94=b+g ??

OpenStudy (jamesj):

Yes. Now, what about a second equation.

OpenStudy (jamesj):

"number of boys is 22 more than twice the number of girls"

OpenStudy (anonymous):

b+22=g

OpenStudy (jamesj):

Not quite. B = G + 22, because the number of boys, B, is more than G.

OpenStudy (jamesj):

Hence you have two equations B + G = 94 B - G = 22 (where I rewrote the second equation, B = G + 22) So far so good?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

why is it b-g=22 ??

OpenStudy (jamesj):

Do you agree that B = G + 22 ?

OpenStudy (jamesj):

If so, then subtract G from each side: B - G = G + 22 - G i.e., B - G = 22

OpenStudy (anonymous):

im confused a little , sorry

OpenStudy (jamesj):

Never mind about that then. What are the two equations from your point of view?

OpenStudy (anonymous):

94=b+g B = G + 22

OpenStudy (jamesj):

Right. B + G = 94 B = G + 22 Two linear equations in two unknowns. Now solve them. Do you know how?

OpenStudy (anonymous):

we wanna get b by it self so - g on each side right ?

OpenStudy (jamesj):

Well, I would substitute the second equation into the first equation.

OpenStudy (jamesj):

Substitute B = G + 22 into B + G = 94

OpenStudy (anonymous):

ok then what do i do my teacher didnt show me how to do thaat ?

OpenStudy (anonymous):

i worked it out and got 24 am i right ??

OpenStudy (jamesj):

This is the standard way to solve simultaneous equations in high school. If you do that then we have B + G = 94 (G + 22) + G = 94 2G + 22 = 94 Now see what to do?

OpenStudy (jamesj):

(and btw, 24 is not the value of B or G)

OpenStudy (anonymous):

yes then you divide 2 into 94 and 2 right ?

OpenStudy (jamesj):

You could divide the entire equation by 2, in which case you would have 2G + 22 = 94 => G + 11 = 47

OpenStudy (anonymous):

yes then u subtract 11 from 47

OpenStudy (jamesj):

Hence G = ... what?

OpenStudy (anonymous):

36 but thats not a choice tho

OpenStudy (jamesj):

For B or G? If G = 36, what does B have to be?

OpenStudy (anonymous):

58?

OpenStudy (jamesj):

Yes B = 58 G = 36 We can verify that is the answer by making sure those values satisfy the equations B + G = 94 B = G + 22 which indeed they do.

OpenStudy (anonymous):

but thats not a choice tho and he 4 answers

OpenStudy (jamesj):

No, it's not. So what's wrong? What's wrong is our second equation didn't correspond to the question: "boys is 22 more than twice the number of girls"

OpenStudy (jamesj):

Hence the second equation should have been B = 2G + 22

OpenStudy (jamesj):

So instead, solve the system B + G = 94 B = 2G + 22

OpenStudy (anonymous):

ohhh meaning it would of been 48

OpenStudy (jamesj):

Really? How'd you get that?

OpenStudy (jamesj):

Work through the algebra again, now with the correct equations.

OpenStudy (anonymous):

bec ause if u divide 2 into 22 it would be 11 then you add that to the to 37 (hence like we did before cus we added the 22 to 37) IN CONCLUSION IT WOULD ADD TO 48!!!! YAY !!!

OpenStudy (jamesj):

Wrong.

OpenStudy (jamesj):

B + G = 94 B = 2G + 22 Substitute the second equation into the first equation.

OpenStudy (anonymous):

IM CONFUSED UGHH

OpenStudy (jamesj):

These are the correct two equations B + G = 94 B = 2G + 22 You need to solve them. The way to solve them is to use the substation technique we saw above. To do that, substitute the second equation into the first equation and eliminate the variable B:

OpenStudy (jamesj):

*substitution

OpenStudy (anonymous):

24

OpenStudy (jamesj):

What's 24?